# Optimal Computation in Leaderless and Multi-Leader Disconnected Anonymous Dynamic Networks

### 20 Feb 2024 v5

@@ Abstract
While most of the existing literature on anonymous dynamic networks relies on classical mass-distribution techniques, our work makes use of a recently introduced combinatorial structure called history tree, also developing its theory in new directions. Among other contributions, our results make definitive progress on two popular fundamental problems for anonymous dynamic networks: leaderless Average Consensus (i.e., computing the mean value of input numbers distributed among the processes) and multi-leader Counting (i.e., determining the exact number of processes in the network). In fact, our approach unifies and improves upon several independent lines of research on anonymous networks, including Nedić et al., IEEE Trans. Automat. Contr. 2009; Olshevsky, SIAM J. Control Optim. 2017; Kowalski–Mosteiro, ICALP 2019, SPAA 2021; Di Luna–Viglietta, FOCS 2022.
@@ 1 Introduction
Dynamic networks. An increasingly prominent area of distributed computing focuses on algorithmic aspects of dynamic networks, motivated by novel technologies such as wireless sensors networks, software-defined networks, networks of smart devices, and other networks with a continuously changing topology [10, 38, 41]. Typically, a network is modeled by a system of $n$ processes that communicate in synchronous rounds; at each round, the network’s topology changes unpredictably.
@@
Disconnected networks. In the dynamic setting, a common assumption is that the network is 1-interval-connected, i.e., connected at all rounds [36, 44]. However, this is not a suitable model for many real systems, due to the very nature of dynamic entities (think of P2P networks of smart devices moving unpredictably) or due to transient communication failures, which may compromise the network’s connectivity. A weaker assumption is that the union of all the network’s links across any $T$ consecutive rounds induces a connected graph on the processes [34, 46]. We say that such a network is $T$-union-connected, and we call $T\geq 1$ its dynamic disconnectivity. [^1] We use the term “disconnected” to refer to $T$-union-connected networks in the sense that they may not be connected at any round. It is worth noting that non-trivial (terminating) computation requires some conditions on temporal connectivity to be met, such as a finite dynamic disconnectivity and its knowledge by all processes (refer to Proposition 2.2).
Anonymous processes. Several works have focused on processes with unique IDs, which allow for efficient algorithms for many different tasks [9, 35, 36, 37, 41, 44]. However, unique IDs may not be available due to operational limitations [44] or to protect user privacy: A famous example are COVID-19 tracking apps, where assigning temporary random IDs to users was not enough to eliminate privacy concerns [50]. Systems where processes are indistinguishable are called anonymous. The study of static anonymous networks has a long history, as well [7, 8, 12, 13, 14, 27, 49, 52].
@@
Networks with leaders. It is known that several fundamental problems for anonymous networks (a notable example being the Counting problem, i.e., determining the total number of processes $n$) cannot be solved without additional “symmetry-breaking” assumptions. The most typical choice is the presence of a single distinguished process called leader [2, 3, 4, 5, 23, 27, 29, 31, 40, 48, 53] or, less commonly, a subset of several leaders (and knowledge of their number [^2] It is easy to see that a network with an unknown number of leaders is equivalent to a network with no leaders at all. Also, if the leaders are distinguishable from each other, then any one of them can be elected as a unique leader. Hence, the only genuinely interesting multi-leader case is the one with a known number of indistinguishable leaders.[30, 32, 33, 34].
@@
Apart from the theoretical importance of generalizing the usual single-leader scenario, studying networks with multiple leaders also has a practical impact in terms of privacy. Indeed, while the communications of a single leader can be traced, the addition of more leaders provides differential privacy for each of them.
@@
Leaderless networks. In some networks, the presence of reliable leaders may not always be guaranteed: For example, in a mobile sensor network deployed by an aircraft, the leaders may be destroyed as a result of a bad landing; also, the leaders may malfunction during the system’s lifetime. This justifies the extensive existing literature on networks with no leaders [16, 17, 42, 43, 45, 51, 54]. Notably, a large portion of works on leaderless networks have focused on the Average Consensus problem, where the goal is to compute the mean of a list of numbers distributed among the processes [6, 15, 16, 26, 46, 47].
@@ 1.1 Our Contributions
Summary. Focusing on anonymous dynamic networks, in this paper we completely elucidate the relationship between leaderless networks and networks with (multiple) leaders, as well as the impact of the dynamic disconnectivity $T$ on the efficiency of distributed algorithms. We remark that only a minority of existing works consider networks that are not necessarily connected at all times.
@@
Computability. We give an exact characterization of which functions can be computed in anonymous dynamic networks with and without leaders, respectively. Namely, with at least one leader, all the so-called multi-aggregate functions are computable; [^3] Another way of stating this result is that it is sufficient to know the size of any subset of distinguished processes in order to compute all multi-aggregate functions. with no leaders, only the frequency-based multi-aggregate functions are computable [^4] A similar result, limited to static leaderless networks, was obtained in [28]. (see Section 2 for definitions). Interestingly, computability is independent of the dynamic disconnectivity $T$. Our contribution considerably generalizes a recent result on the functions computable with exactly one leader and with $T=1$ [24].
Complete problems. While computing the so-called Generalized Counting function $F_{GC}$ was already known to be a complete problem for the class of multi-aggregate functions [24], in this work we expand the picture by identifying a complete problem for the class of frequency-based multi-aggregate functions, as well: the Frequency function $F_{R}$ (both $F_{GC}$ and $F_{R}$ are defined in Section 2). By “complete problem” we mean that computing such a function allows the immediate computation of any other function in the class with no overhead in terms of communication rounds.
@@
Algorithms. We give efficient deterministic algorithms [^5] An advantage of using randomization would be the possibility of choosing unique IDs with high probability, but this would not achieve an asymptotic improvement on the running times of our algorithms; see [36]. for computing the Frequency function (Section 3) and the Generalized Counting function (Section 4). Since the two problems are complete, we automatically obtain efficient algorithms for computing all functions in the respective classes.
@@
For each problem, we give two algorithms: a terminating version, where each process is required to commit on its output and never change it, and a stabilizing version, where processes are allowed to modify their outputs, provided that they eventually stabilize on the correct output.
@@
The stabilizing algorithms for both problems run in $2Tn$ rounds regardless of the number of leaders, and do not require any knowledge of the dynamic disconnectivity $T$ or the number of processes $n$. Our terminating algorithm for leaderless networks runs in $T(n+N)$ rounds with knowledge of $T$ and an upper bound $N\geq n$; the terminating algorithm for $\ell\geq 1$ leaders runs in $(\ell^{2}+\ell+1)Tn$ rounds [^6] Note that the case where all processes are leaders is not equivalent to the case with no leaders, because processes do not have the information that $\ell=n$, and have to “discover” that there are no non-leader processes in the network. with no knowledge about $n$. The latter running time is reasonable (i.e., linear) in most applications, as $\ell$ is typically a constant or very small compared to $n$.
@@
Negative results. Some of our algorithms assume processes to have a-priori knowledge of some parameters of the network; in Section 5 we show that all of these assumptions are necessary. We also provide lower bounds that asymptotically match our algorithms’ running times, assuming that the number of leaders $\ell$ is constant (which is a realistic assumption in most applications).
@@
Multigraphs. All of our results hold more generally if networks are modeled as multigraphs, as opposed to the simple graphs traditionally encountered in nearly all of the literature. This is relevant in many applications: in radio communication, for instance, multiple links between processes naturally appear due to the multi-path propagation of radio waves.
1.2 Technical Advances
Our approach departs radically from the mass-distribution techniques traditionally adopted by most previous works on anonymous dynamic networks [18, 19, 20, 21, 32, 33, 34]; instead, we build upon history trees, a combinatorial structure recently introduced in [24]. Each node in a history tree represents a set of processes that are “indistinguishable” at a certain point in time; the number of such processes is the anonymity of the node (see Section 2 for a proper definition of history tree). The theory presented in [24] leads to an optimal solution to the Generalized Counting problem for always connected networks with a unique leader. In this paper we extend the theory to leaderless and multi-leader disconnected dynamic networks thanks to the following technical breakthroughs.
  • Main contribution. We succeed in generalizing the theory in [24] to networks with more than one leader (Section 4.2). This is an especially challenging task because there is no obvious way of adapting the counting algorithm from [24], which critically assumes the existence of a single node in the history tree representing the leader, as well as the exact knowledge of its anonymity. In contrast, in the history tree of a multi-leader network there may be multiple nodes representing leaders, and the only information available is that the sum of their anonymities is $\ell$.
    To work around this difficulty, we develop an approximate counting procedure which repeatedly makes “hypotheses” on the anonymity of a selected leader node. Starting from a given hypothesis, we show how to derive conditional anonymities for other nodes in the history tree, until we are able to estimate the network’s size. Furthermore, we show that the correctness of an estimate can be verified efficiently; if the estimate is incorrect, we proceed with a new hypothesis on the anonymity of a leader’s node, and so on. If carefully implemented, this strategy leads to a correct counting algorithm for multi-leader networks (Theorem 4.3).
  • Secondary contribution. We introduce a novel technique to transform a history tree into a system of independent linear equations on anonymities (Lemma 3.1). This technique is the basis for a variety of optimal algorithms: It yields a stabilizing algorithm for the Frequency function in leaderless networks (Section 3.1), which can be converted into a terminating algorithm if a bound on the propagation time of information is known (Section 3.2), as well as a stabilizing counting algorithm for multi-leader networks (Section 4.1).
  • Additional contribution. All of our algorithms and techniques apply to multigraphs, i.e., networks that may have multiple parallel links between pairs of processes. [^7] The original design of history trees is general enough to be applied to networks modeled as multigraphs [24]. However, this fact alone does not automatically imply that any algorithm that utilizes history trees should work for any network that can be modeled as a multigraph. There are tasks that can only be carried out on networks with no parallel links which can still benefit from a formulation in terms of history trees. In fact, a history tree is just an alternative way of representing a network, and making good use of this representation is entirely up to the algorithm. This turns out to be a remarkably powerful feature in light of Proposition 2.3, which establishes a relationship between multigraphs and $T$-union-connected networks. This finding single-handedly allows us to generalize our algorithms to disconnected networks at the cost of a mere factor of $T$ in their running times, which is worst-case optimal. This significantly improves upon previous counting algorithms for disconnected networks, which had an exponential dependence on $T$ (see, e.g., [33]).
1.3 Impact on Fundamental Problems and State of the Art
As a byproduct of the results mentioned in Section 1.1, we are able to optimally solve two popular fundamental problems: Generalized Counting for multi-leader networks (because it is a multi-aggregate function) and Average Consensus for leaderless networks (because the mean is a frequency-based multi-aggregate function). As summarized in Table 1 and as discussed below, our results improve upon the state of the art on both problems in terms of (i) running time, (ii) assumptions on the network and the processes’ knowledge, and (iii) quality of the solution. Altogether, we settle open problems from ICALP 2019 [30], SPAA 2021 [32], and FOCS 2022 [24]. For a more thorough discussion and a comprehensive survey of related literature, refer to Appendix B.
Average Consensus. This problem has been studied for decades by the distributed control and distributed computing communities [6, 15, 16, 17, 32, 42, 45, 47, 51, 54]. In the following, we argue that our results directly improve upon the current state of the art on this problem. A more detailed discussion can be found in the surveys [26, 43, 46] and in Appendix B.
A convergent algorithm with a running time of $O\left(Tn^{3}\log(1/{\epsilon})\right)$ is given in [42]. The algorithm works in $T$-union-connected networks with no knowledge of $T$, but it rests on the assumption that the degree of each process in the network has a known upper bound. Assuming an always connected network, [16] gives an algorithm that converges in $O\left(n^{4}\log(n/{\epsilon})\right)$ rounds. We remark that both algorithms are only $\epsilon$-convergent; therefore, not only does our stabilizing algorithm improve upon their running times, but it solves a more difficult problem under weaker assumptions. [^8] More generally, averaging algorithms based on Metropolis rules cannot be applied to our model, because they require all processes to know their out-degree before the broadcast phase of each round. This was noted by Charron-Bost et al. in [16]: “Unfortunately, local algorithms cannot implement the Metropolis rule over dynamic networks. […] an agent’s next estimate $x_{i}(t)$ depends on information present within distance $2$ of agent $i$ in the communication graph $\mathbb{G}(t)$, which is not local enough when the network is subject to change”.
The algorithm in [15] stabilizes to the actual average in a linear number of rounds, but it is a randomized Monte Carlo algorithm and requires the network to be connected at each round. In contrast, our linear-time stabilizing algorithm is deterministic and works in disconnected networks.
As for terminating algorithms, the one in [32] terminates in $O\left(n^{5}\log^{3}({n})/\ell\right)$ rounds assuming the presence of a known number $\ell$ of leaders and an always connected network. Since the number of leaders is known, our terminating algorithm for Generalized Counting also solves Average Consensus with a running time that improves upon [32] and does not require the network to be connected. We remark that our algorithm terminates in linear time when $\ell$ is constant.
Generalized Counting. Our results on this problem are direct generalizations of [24] to the case of multiple leaders and disconnected networks. The best previous counting algorithm with multiple known leaders is the one in [34], which terminates in $O\left({n^{4}}\log^{3}(n)/\ell\right)$ rounds and assumes the network to be connected at each round. In the same setting, our stabilizing and terminating algorithms have running times of $2n$ rounds and $(\ell^{2}+\ell+1)n$ rounds, respectively.
The only other result for disconnected networks is the recent preprint [33], which gives an algorithm that terminates in $\widetilde{O}\left({n^{2T+3}}/{\ell}\right)$ rounds using $O(\log n)$-sized messages. Our terminating algorithm has a linear dependence on both $n$ and $T$, which is an exponential improvement upon the running time of [33], but it requires polynomial-sized messages.
Problem Reference Leaders Disconn. Term.
Notes
Running time
Average Consensus [42] $\ell=0$
$\epsilon$-convergence, $T$ unknown, upper bound on processes’ degrees known
$O(Tn^{3}\log({1}/{\epsilon}))$
[16] $\ell=0$
$\epsilon$-convergence
$O(n^{4}\log({n}/{\epsilon}))$
[15] $\ell=0$
randomized Monte Carlo
$O(n)$
[32] $\ell\geq 1$
$\ell$ known
$O({n^{5}}\log^{3}({n})/\ell)$
this work $\ell=0$
$T$ unknown
$2Tn$
$\ell=0$
$T$ and $N\geq n$ known
$T(n+N)$
(Generalized) Counting [24] $\ell=1$ $22n-$ 2
$\ell=1$ $23n-$ 2
[34] $\ell\geq 1$
$\ell$ known
$O({n^{4}}\log^{3}(n)/\ell)$
[33] $\ell\geq 1$
$\ell$ and $T$ known, $O(\log n)$-size messages
$\widetilde{O}({n^{2T+3}}/{\ell})$
this work $\ell\geq 1$
$\ell$ known, $T$ unknown
$2Tn$
$\ell\geq 1$
$\ell$ and $T$ known
$(\ell^{2}+\ell+1)Tn$
Table 1: Comparing results for Average Consensus and Counting in anonymous dynamic networks. For algorithms that support disconnected networks, $T$ indicates the dynamic disconnectivity.

### 05 Apr 2024 v6

# Optimal Computation in Anonymous Dynamic Networks

@@ Abstract
While most of the existing literature on anonymous dynamic networks relies on classic mass-distribution techniques, our work makes use of a novel combinatorial structure called history tree, which is of independent interest. Among other contributions, our results make conclusive progress on two popular fundamental problems for anonymous dynamic networks: leaderless Average Consensus (i.e., computing the mean value of input numbers distributed among the processes) and multi-leader Counting (i.e., determining the exact number of processes in the network). In fact, our approach unifies and improves upon several independent lines of research on anonymous networks, including Yamashita–Kameda, IEEE T. Parall. Distr. 1996; Boldi–Vigna, Discrete Math. 2002; Nedić et al., IEEE T. Automat. Contr. 2009; Kowalski–Mosteiro, J. ACM 2020.
Our contribution not only opens a promising line of research on applications of history trees, but also demonstrates that computation in anonymous dynamic networks is practically feasible and far less demanding than previously conjectured.
@@ 1 Introduction
1.1 Background
Dynamic networks. The study of theoretical and practical aspects of highly dynamic distributed systems has garnered significant attention in recent years [11, 40, 43]. These systems involve a constantly changing network of computational devices called processes (sometimes referred to as “processors” or “agents”). Pairs of processes may send each other messages only through links that appear or disappear unpredictably. This dynamicity is typical of modern real-world systems and is the result of technological innovations, such as the spread of mobile devices, software-defined networks, wirelesses sensor networks, wearable devices, smartphones, etc.
Connected networks. There are several models of dynamicity [11]; a popular choice is the 1-interval-connected network model [38, 47]. Here, a fixed set of $n$ processes communicate through links forming a time-varying graph, i.e., a graph whose edge set changes at discrete time units called rounds (thus, the system is synchronous); such a graph changes unpredictably, but is assumed to be connected at all times.
@@ 1.1 Background
Disconnected networks. The 1-interval-connected network model may not be a suitable choice for many real systems, due to the very nature of dynamic entities (think of P2P networks of smart devices moving unpredictably) or due to transient communication failures, which may compromise the network’s connectivity. A more realistic assumption is that the union of all the network’s links across any $\tau$ consecutive rounds induces a connected graph on the processes [36, 44, 49]. [^1] It is well known that non-trivial (terminating) computation requires some conditions on temporal connectivity to be met, as well as some a-priori knowledge by processes (refer to Proposition 2.3). We say that such a network is $\tau$-union-connected, and the smallest such parameter $\tau\geq 1$ is called dynamic disconnectivity. [^2] Another widely used parameter for dynamic networks is the dynamic diameter $d$. The relationship between $\tau$ and $d$ is discussed at the end of Section 2.3. In particular, it is worth noting that every occurrence of the parameter $\tau$ in the running times of our algorithms can be safely replaced with $d$. Observe that 1-interval-connected networks can equivalently be characterized as 1-union-connected networks.
Networks with unique IDs. A large number of research papers have considered dynamic systems where each process has a distinct identity (unique IDs[38]. In this setting, there are efficient algorithms for consensus [39], broadcast [10], counting [38, 47], and many other fundamental problems [37, 43].
It should be noted that networks with unique IDs allow for very simple algorithms for a wide variety of problems. Indeed, in a 1-interval-connected network of $n$ processes, if each process broadcasts the set of its “known IDs” at every round (initially, a process only knows its own ID), then every process can learn the IDs of all other processes in $n$ rounds. If input values are attached to these IDs, then the network can compute any function of such input values within $n$ rounds.
Anonymous networks. Networks without IDs, also called anonymous systems, are a considerably more challenging scenario. In this model, all processes have identical initial states, and may only differ by their inputs. In the last thirty years, a large body of works have investigated the computational power of anonymous static networks, giving characterizations of what can be computed in various settings [8, 9, 13, 14, 15, 28, 52, 56].
Anonymous systems are not only important from a theoretical standpoint, but also have a remarkable practical relevance. In a highly dynamic system, IDs may not be guaranteed to be unique due to operational limitations [47], or may compromise user privacy. Indeed, users may not be willing to be tracked or to disclose information about their behavior; examples are COVID-19 tracking apps [53], where a threat to privacy was felt by a large share of the public even if these apps were assigning a rotating random ID to each user. In fact, an adversary can easily track the continuous broadcast of a fixed random ID tracing the movements of a person [4]. Anonymity is also found in insect colonies and other biological systems [29].
@@
Networks with leaders. In order to deterministically solve several non-trivial problems in anonymous systems, it is necessary to have some form of initial “asymmetry” [1, 9, 42, 56]. The most common assumption is the existence of a single distinguished process called leader [2, 3, 5, 6, 22, 28, 31, 33, 42, 51, 57] or, less commonly, a subset of several leaders (and knowledge of their number [^3] It is easy to see that a network with an unknown number of leaders is equivalent to a network with no leaders at all. Also, if the leaders are distinguishable from each other, then any one of them can be elected as a unique leader. Hence, the only genuinely interesting multi-leader case is the one with a known number of indistinguishable leaders.[32, 34, 35, 36]. The presence of leaders is a realistic assumption: examples include base stations in mobile networks, gateways in sensor networks, etc. For these reasons, the computational power of anonymous systems enriched with one or more leaders has been extensively studied in the traditional model of static networks [28, 51, 57], as well as in population protocols [2, 3, 5, 6, 22].
@@
Apart from the theoretical importance of generalizing the usual single-leader scenario, studying networks with multiple leaders also has practical impacts in terms of privacy. Indeed, while the communications of a single leader can be traced, the addition of more leaders provides differential privacy for each of them.
@@
Leaderless networks. In some networks, the presence of reliable leaders may not always be guaranteed. For example, in a mobile sensor network deployed by an aircraft, the leaders may be destroyed as a result of a bad landing; also, the leaders may malfunction during the system’s lifetime. This justifies the extensive existing literature on networks with no leaders [16, 44, 45, 48, 55, 58]. Notably, a large portion of works on leaderless networks have focused on the Average Consensus problem, where the goal is to compute the mean of a list of numbers distributed among the processes [7, 26, 49, 50].
Problem classes. Assume that each process is assigned an input at the beginning of the computation, i.e., at round 00. The Input Frequency function is the function that returns the percentage of processes that have each input value. Being able to compute the Input Frequency function allows a system to compute a wide class of functions called frequency-based with no loss in performance [30]. Thus, we say that the Input Frequency function is complete for this class of functions (refer to Section 2.2). The most prominent representative of this class of functions is given by the aforementioned Average Consensus problem, since the mean of a multiset of numeric input values is a frequency-based function (the percentage of processes that have each input can be used as weight to compute the mean of all inputs).
The Input Multiset function is the function that returns the number of processes that have each input value. This function is complete for a class of functions called multiset-based, which strictly includes the frequency-based ones (see Section 2.2 for definitions). A well-studied example of a multiset-based function that is not frequency-based is given by the Counting problem, which asks to determine the total number of processes in the system. The Counting problem is practically interesting in real-world scenarios such as large-scale ad-hoc sensor networks, where the individual processes may be unaware of the size of the system [47].
1.2 State of the Art
Average Consensus problem. In the Average Consensus problem, each process starts with an input value, and the goal is to compute the average of these initial values. The typical approach found in the literature is based on local averaging algorithms, where each process updates its local value at each round based on a convex combination of the values of its neighbors. This and similar techniques for Average Consensus have been studied for decades by the distributed control and distributed computing communities  [7, 16, 34, 44, 48, 49, 50, 55, 58]. Existing works can be divided into those that give convergent solutions, those that give stabilizing solutions, and those that give terminating solutions.
In convergent algorithms, the consensus is not reached in finite time, but each process’s local value asymptotically converges to the average. A state-of-the-art $\epsilon$-convergent algorithm based on Metropolis rules with a running time of $O\left(\tau n^{3}\log(1/{\epsilon})\right)$ communication rounds is given in [44]. However, this algorithm rests on the assumption that the degree of each process in the network has a known upper bound. [^4] More generally, averaging algorithms based on Metropolis rules cannot be applied to our network model, because they require all processes to know their outdegree at every round prior to sending their messages. The running time can be improved to $O\left(n^{2}\log(1/\epsilon)\right)$ communication rounds if the network is always connected (i.e., $\tau=1$) and may only change every three rounds (as opposed to every round).
In stabilizing algorithms, the consensus is reached in a finite number of rounds, but no termination criteria are specified, and therefore all processes are assumed to run indefinitely. To the best of our knowledge, all stabilizing Average Consensus algorithms for leaderless networks are either probabilistic or assume the network to be static [26, 30, 48, 58]. However, some of these algorithms stabilize in a linear number of communication rounds [48, 58].
As for terminating Average Consensus algorithms, the one in [34] terminates in $O\left(n^{5+\epsilon}\log^{3}({n})/\ell\right)$ communication rounds (for $\epsilon>0$) assuming the presence of a known number $\ell\geq 1$ of leaders and an always connected network.
A major research question left open in previous works is the following.
Research Question 1. In anonymous dynamic leaderless networks, are there deterministic Average Consensus algorithms that stabilize (or terminate) in linear time?
Counting problem. A long series of papers have focused the Counting problem in connected anonymous networks with a unique leader [12, 17, 18, 19, 20, 21, 31, 32, 36, 42]. [^5] It is folklore that the Counting problem cannot be solved in leaderless networks. These works have achieved better and better running times, with the first polynomial-time algorithm presented in J. ACM in 2020, having a running time of $O\left({n^{5}}\log^{2}(n)\right)$ rounds [33]. This was later improved and extended to networks with $\ell\geq 1$ leaders in [36], solving the Counting problem in $O\left({n^{4+\epsilon}}\log^{3}(n)/\ell\right)$ communication rounds (for $\epsilon>0$).
The only result for $\tau$-union-connected networks is the recent preprint [35], which gives an algorithm that terminates in $\widetilde{O}\left({n^{3+2\tau(1+\epsilon)}}/{\ell}\right)$ communication rounds, assuming the presence of $\ell\geq 1$ leaders. We remark that this algorithm is exponential in $\tau$, and for $\tau=1$ it has a running time of $\widetilde{O}\left({n^{5+\epsilon}}/{\ell}\right)$.
Almost all of these works share the same basic approach of implementing a mass-distribution mechanism similar to the local averaging used to solve the Average Consensus problem. [^6] We remark that the mass-distribution approach used for Counting and Average Consensus requires processes to communicate numbers whose representation size grows at least linearly with the number of rounds. Thus, all cited algorithms require processes to exchange messages of polynomial size. In spite of the technical sophistication of this line of research, there is still a striking gap in terms of running time between state-of-the-art algorithms for anonymous networks and networks with unique IDs. The same gap exists with respect to static anonymous networks, where the Counting problem is known to be solvable in $O(n)$ rounds [42]. Given the current state of the art, solving non-trivial problems in large-scale dynamic networks is still impractical.
Research Question 2. In anonymous dynamic networks with a fixed number of leaders, are there deterministic Counting algorithms that stabilize (or terminate) in linear time?
@@ 1.3 Our Contributions
Problem class Leaders Terminating Assumptions Running time Lower bound
Frequency-based (e.g., Average Consensus) $\ell=0$ $\tau(2n-2)+1$ $\tau(2n-6)$
$\tau$ and $N\geq n$ known $\tau(n+N-2)+1$ $\tau(2n-4)$
$d$ known $\tau(n-1)+d+1\leq dn+1$
[…]
$\ell$ and $\tau$ known $\tau((\ell^{2}+\ell+1)(n-1)+1)$ $\tau(2n-\ell-3)$
Table 1: Summary of the results in this paper. The variable $n$ indicates the number of processes in the system, $\ell$ is the number of leaders, $N$ is an upper bound on $n$, $\tau$ is the dynamic disconnectivity of the network, and $d$ is (an upper bound on) its dynamic diameter. If $\ell=0$, only the frequency-based functions are computable; if $\ell\geq 1$, the multiset-based functions are also computable.
@@
Summary. Focusing on anonymous dynamic networks, in this paper we completely elucidate the relationship between leaderless networks and networks with one or more leaders, as well as the impact of the dynamic disconnectivity $\tau$ on the efficiency of distributed algorithms. In particular, we characterize the solvable problems in each of these settings and we provide optimal linear-time algorithms for all solvable problems.
Technique. Our approach departs radically from the mass-distribution and averaging techniques traditionally adopted by most previous works on anonymous dynamic networks. Instead, all of our results are based on a novel combinatorial structure called history tree, which completely represents an anonymous dynamic network and naturally models the idea that processes can be distinguished if and only if they have different “histories” (see Section 3). Thanks to the simplicity of our technique, this paper is entirely self-contained, our proofs are transparent and easy to understand, and our algorithms are elegant and straightforward to implement.
An implementation can be found at https://github.com/viglietta/Dynamic-Networks. The repository includes a dynamic network simulator that can be used to test our algorithms and visualize the history trees of custom networks.
@@
Computability. We give an exact characterization of the set of functions that can be computed deterministically in anonymous dynamic networks with and without leaders, respectively. Namely, in networks with at least one leader, a function is computable if and only if it is multiset-based; [^7] Another way of stating this result is that it is sufficient to know the size of any subset of distinguished processes in order to compute all multiset-based functions. in networks with no leaders, a function is computable if and only if it is frequency-based. [^8] A similar result, limited to static leaderless networks and functions with additional constraints, was obtained in [30]. Interestingly, computability is independent of the dynamic disconnectivity $\tau$.
@@
Algorithms. Furthermore, we give efficient deterministic [^9] An advantage of allowing randomization in our model would be the possibility of choosing unique IDs with high probability. Even though this would greatly simplify most problems, it would not afford an asymptotic improvement on the running times of our algorithms; see [38]. algorithms for computing the Input Frequency function in leaderless networks (Sections 4.2 and 4.3) and the Input Multiset function in networks with leaders (Sections 4.4 and 5). Since these functions are complete for the class of frequency-based and multiset-based functions, respectively, we automatically obtain efficient algorithms for computing all functions in these classes (Section 2.2).
@@
For each of the aforementioned functions, we give two algorithms: a terminating version, where all processes are required to commit on their output and never change it, and a slightly more efficient stabilizing version, where processes are allowed to modify their outputs, provided that they eventually stabilize on the correct output. The running times of our algorithms are summarized in Table 1.
@@
The stabilizing algorithms for both functions give the correct output within $2\tau n$ communication rounds regardless of the number of leaders, and do not require any knowledge of the dynamic disconnectivity $\tau$ or the number of processes $n$. Our terminating algorithm for leaderless networks runs in $\tau(n+N)$ communication rounds with knowledge of $\tau$ and an upper bound $N\geq n$; the terminating algorithm for $\ell\geq 1$ leaders runs in $\tau(\ell^{2}+\ell+1)n$ communication rounds [^10] Note that the case where all processes are leaders is not equivalent to the case with no leaders, because processes do not have the information that $\ell=n$, and have to “discover” that there are no non-leader processes in the network. with no knowledge of $n$. The latter running time is reasonable (i.e., linear) in most applications, as $\ell$ is typically a constant or negligible compared to $n$.
We emphasize that the running times of our algorithms are specified as precise values, rather than being expressed in big-O notation. In fact, we made an effort to optimize the multiplicative constants, as well as the asymptotic complexity of our algorithms. To our knowledge, this feature is unique within the entire body of literature on anonymous dynamic networks.
We remark that the local computation time and the amount of processes’ internal memory required by our terminating algorithms is only polynomial in the size of the network. Also, like in previous works on anonymous dynamic networks, processes need to send messages of polynomial size.
@@
Negative results. Some of our algorithms assume processes to have a-priori knowledge of some parameters of the network; in Section 6 we show that none of these assumptions can be removed. We also provide lower bounds that asymptotically match our algorithms’ running times, assuming that the number of leaders $\ell$ is constant (which is a realistic assumption in most applications). Table 1 summarizes our lower bounds, as well. We point out that our lower bounds of roughly $2\tau n$ rounds are the first non-trivial lower bounds for simple undirected anonymous dynamic networks (i.e., better than $n-1$).
@@
Multigraphs. All of our results hold more generally if networks are modeled as multigraphs, as opposed to the simple graphs traditionally encountered in nearly all of the existing literature. This is relevant in many applications: in radio communication, for instance, multiple links between processes naturally appear due to the multi-path propagation of radio waves. Furthermore, this turns out to be a remarkably powerful feature in light of Proposition 2.4, which establishes a relationship between multigraphs and $\tau$-union-connected networks. This finding single-handedly allows us to generalize our algorithms to disconnected networks at the cost of a mere factor of $\tau$ in their running times, which is worst-case optimal.
Significance. Our general technique based on history trees enables us to approach all problems related to anonymous networks in a uniform and systematic manner. This technique also allows us to design straightforward algorithms that achieve previously unattainable running times. Indeed, our results make conclusive advancements on long-standing problems within anonymous dynamic networks, particularly with regards to the Average Consensus and Counting problems, in several aspects:
  • Running time. Our algorithms are asymptotically worst-case optimal, i.e., linear (for networks with a fixed number of leaders $\ell$ and a fixed dynamic disconnectivity $\tau$).
  • Assumptions on the network. Unlike most previous solutions, our algorithms work in any dynamic network that has a finite dynamic disconnectivity.
  • Knowledge of the processes. Our algorithms assume that the processes have a-priori knowledge about certain properties of the network only when the absence of such knowledge would render the Average Consensus or Counting problems unsolvable.
  • Quality of the solution. Unlike several previous works, our algorithms are deterministic rather than probabilistic, and stabilize or terminate rather than converge.
Altogether, we settle open problems from several papers, including Nedić et al., IEEE Trans. Automat. Contr. 2009 [44]; Olshevsky, SIAM J. Control Optim. 2017 [48]; Kowalski–Mosteiro, J. ACM 2020 [33] and SPAA 2021 [34]. In particular, we settle the two research questions stated in Section 1.2:
Contribution 1. In anonymous dynamic leaderless networks, any problem that can be solved deterministically (including Average Consensus) has an algorithm that terminates (thus stabilizes) in linear time.
Contribution 2. In anonymous dynamic networks with a fixed number of leaders, any problem that can be solved deterministically (including Counting) has an algorithm that terminates (thus stabilizes) in linear time.
Our findings indicate that computing within anonymous dynamic networks entails an overhead when compared to networks equipped with unique IDs. However, this overhead is only linear, which is substantially lower than previously conjectured [33, 36]. In fact, our results demonstrate that general computations in anonymous and dynamic large-scale networks are both feasible and efficient in practice.
Unique-leader case. Our results concerning connected networks with a unique leader (i.e., $\tau=\ell=1$) are especially noteworthy. We have an algorithm for the Counting problem that stabilizes in less than $2n$ rounds, which is worst-case optimal up to a small additive constant. We also have a terminating algorithm with a running time of less than $3n$ rounds, which is remarkably close to the lower bound of $2n$ rounds. We recall that, prior to our work, the state-of-the-art algorithm for this problem had a running time of $O\left({n^{4+\epsilon}}\log^{3}(n)\right)$ rounds (with $\epsilon>0$[36].
Previous versions. This work is an extension of two preliminary conference papers that appeared at FOCS 2022 [23] and DISC 2023 [25], respectively. In preparing the current version, we have reworked several sections, included all missing proofs, and provided a more accurate analysis of the performance of our algorithms. In addition, the upper bound in terms of the dynamic diameter $d$ for leaderless networks is a new contribution (see Table 1). We also added a comparisons between history trees and previous structures and concepts related to anonymous and dynamic networks (see Section 3.3).